a bicycle coasting at 9.0 m/s comes to a 5.0 m long, 1.5m high ramp.

what is the bicycles speed as it leaves the top of the ramp?
Answer #1

D Drawing Condition - initial reloudy = amk V = amis H=1.5m from Conservation of Enfy I mu? = might me ²-2gh va 2 - 2x9.8 x
Answer #2

SOLUTION :


K. E. at the beginning of ramp = 1/2 m v^2 = 1/2 m (9)^2 = 40.5 m   (N-m) 


P. E at the top of the ramp = m g h = m * 9.8 * 1.5 = 14.7 m    (N-m)


Let the speed at the top of ramp be V m/sec.

So K.E at the top of the ramp = 1/2 m V^2 


So, aș per energy conservation :


Total energy at the beginning of the ramp = Total energy at the top of the ramp

=> 40.5 m = 14.7 m + 1/2 m V^2 

=> 40.5 = 14.7 + 1/2 V^2

=> V^2 = (40.5 - 14.7) * 2 = 51.6 

=> V = sqrt(51.6) = 7.18 m/sec 


So, bicycle speed at the top of ramp = 7.18 m/sec (ANSWER).


answered by: Tulsiram Garg